OpenStudy (anonymous):

what is y^2-3y-2/y^2+5y+6 divided by y^2+2y-8/y^2+y-12 ?

OpenStudy (ledah):

\[y^{2}-3y-2/y^{2}+5y+6 \div y^{2}+2y-8/y ^{2}+y-12\]

OpenStudy (ledah):

that it?

OpenStudy (ledah):

your problem I mean

OpenStudy (anonymous):

factor both fractions... then rewrite the division of fractions into multiplying by reciprocals. then cancel common factors.. what u got so far?

OpenStudy (anonymous):

(y-1) / (y+2)(y+3) x (y-3) = (y-1)(y-3) / (y+2)(y+3)

OpenStudy (anonymous):

Thats the answer.

OpenStudy (anonymous):

@Ledah yes that's the proble, @Isaacisbeast123 thanks :D

OpenStudy (jiteshmeghwal9):

\[\frac{y^2-3y-2}{y^2+5y+6} \times \frac{y^2+2y-8}{y^2+y-12}\]